[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Oct 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Oct 7 18:21:02 UTC 2017


HSU Shift Log for Sat Oct 07 2017 14:20:56 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: ASCAT outage-- seems to be NOAA-wide as SDM doesn't  
have any data either. They are working it.

Miscellaneous issues: GOES Meso active until 6Z- evening shift should renew
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171007/9ad2e95d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list